OpenStudy (raffle_snaffle):

Mercury is 13.6 times as dense as liquid water. What would be the reading of a water-filled barometer at normal atmospheric pressure, 760. mmHg?

OpenStudy (raffle_snaffle):

@Abhisar

OpenStudy (abhisar):

What do u think ?

OpenStudy (raffle_snaffle):

I hate just randomly guessing which formula to use.

OpenStudy (raffle_snaffle):

hmmmm

OpenStudy (raffle_snaffle):

Do we use PV=nRT, P being 1 atm, then we can solve for n or find n by looking at the periodic table

OpenStudy (raffle_snaffle):

or do we need an equation with density in it? like for example m/v=MP/RT

OpenStudy (abhisar):

actually m not sure about this one. I think @somy will help u better

OpenStudy (raffle_snaffle):

oh okay.

OpenStudy (aaronq):

were not interested in using the temperature, we're only making a comparison of the volumes between the two liquids at a certain pressure. We can use the definition of density: \(\sf \rho=\dfrac{m}{V}\) The atmospheric pressure will equal masses of both of the liquids up the tube in the barometer (which will give us our reading). Can you think of how to do this now?

OpenStudy (aaronq):

"The atmospheric pressure will push** equal.."

OpenStudy (abmon98):

1 atm=760mmHg=760 tor Three factors affect the pressure in a fluid: The depth the pressure is measured at The density of the fluid The gravitational field strength P=dgh 760=1*10*h h=76

OpenStudy (aaronq):

so you're saying that the reading will be 76 mm? 1 tenth of that of mercury?

OpenStudy (abmon98):

yes

OpenStudy (aaronq):

that's impossible, mercury is much denser than water. The same mass needs to be raised.

OpenStudy (abmon98):

H1/H2=D2/D1 Density of wate is 1.00g/ml Height of H2O/Height of Hg=Density Of Hg/Density of H2O 13.6(g/ml)*760mm/1.00(g/ml)

OpenStudy (aaronq):

yeah, i think that's right.